OpenStudy (anonymous):

If every face of the prism shown below is a square and the volume of the prism is 27cm cubed, what is the surface area?

OpenStudy (anonymous):

where is the prism?

OpenStudy (anonymous):



OpenStudy (anonymous):

ok thanks one sec

OpenStudy (anonymous):

thank you. (:

OpenStudy (anonymous):

A cube with side length s has 6 faces, each with an area of s^2. So SA = 6s^2 Volume = Area of base times height for any right prism, so V = s^2 * s = s^3 SA = 6s^2 = 11 566 s^2 = 1927.666667 s = 43.90520091 V = s^3 = 43.90520091^3 = 84634.59228

OpenStudy (anonymous):

so now reverse that

OpenStudy (anonymous):

how do i reverse that ? because i also have another problem that gives me the volume but its asking m to find the surface area

OpenStudy (anonymous):

ok got it now let me draw

OpenStudy (anonymous):